OpenStudy (anonymous):

an amusement park wants a square pyramid in its ancient wonders theme area. the pyramid is to be 8 ft high and have a 12 ft wide square base. what is the surface area of the pyramid

OpenStudy (lgbasallote):

do you know how to take the surface area of a square prism?

OpenStudy (anonymous):

no did not learn that i have the answer but i just want to check my answers to see if i got the same or if i should try the problem again

OpenStudy (anonymous):

@igbasallote please help

OpenStudy (lgbasallote):

the formula for square prism is 4sh + s^2 h where s is the side of the square and h is the height. now to take the surface area of a square pyramid, just divide the formula by 3 \[SA = \frac{4sh + s^2 h}{3}\] okay so far?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@igbasallote u there?

OpenStudy (lgbasallote):

sorry...that formula is a mistake...i couldn't correct it because i keep lagging up

OpenStudy (anonymous):

i got 336ft squared

OpenStudy (lgbasallote):

how did you get that?

OpenStudy (anonymous):

i did 12 x 12 + 6xthe sh which i got 10 for and then that answer x 4

OpenStudy (lgbasallote):

let me show you the formula first.. formula for surface area of a square prism is \[4sh + 2s^2\] formula for surface area of square pyramid is \[\frac{4sh + 2s^2}3\]

OpenStudy (anonymous):

ok but what do you get as an answer i got 336 and i tryed it 4 times but the book answers says 384

OpenStudy (lgbasallote):

it says side of square is 12 and h is 8 \[\implies \frac{4(12)(8) + 2(12)^2}3\] is that what you did?

OpenStudy (anonymous):

no i dont understand that, my teacher told me to do b x h to find the area of the base and than 1/2 b x h x 4 to find the triangle faces and then add them together

OpenStudy (anonymous):

@igbasallote